Fermi National Laboratory


Accelerator Update for July 22, 2005 - July 25, 2005

Friday July 22
The day shift began with Operations monitoring the Recycler stash, the Tevatron (TeV) recovering from a quench, with the Antiproton Source (Pbar) stacking, with MiniBooNE off due to a breaker repair, and with the NuMI and Meson MCenter experiments taking beam.

MiniBooNE resumed taking beam at 9:48 AM.

The Meson MTest experiment began taking beam at 10:04 AM.

At 10:20, Feeder 47 tripped off and sent a glitch along the power system that tripped off many devices, but the stack and stash were not affected. Feeder 47 had just been reset earlier this morning. It tripped with no load.

Operations established beam to the Booster at 12:04 PM.

Pbar resumed stacking at 12:27 PM.

MiniBooNE resumed taking beam at 12:32 PM.

Operations put the TeV through a wet squeeze at 3:07 PM.

Operations put the TeV into shot setup at 4:19 PM.

Operations established combination store 4293 at 7:17 PM with an initial luminosity of 86.3E30. Pbar resumed stacking.

At 8:44 PM, the Main Injector (MI) tripped off due to MIPSLOOP interlock dropping out. This was due to a MI-10 ground fault indication. It reset okay.

At 9:01 PM, the MI beam permit tripped off due to beam problems. It reset okay.

At 10:36 PM, MI tripped off due to another MI-10 ground fault indication. It reset okay.

Operations put the Recycler into shot setup at 11:15 PM to stash antiprotons.

Saturday July 23
The midnight shift began with Operations monitoring store 4293, with Pbar and the Recycler stashing antiprotons, and with MiniBooNE and NuMI taking beam.

The Recycler stash ended and Pbar resumed stacking at 12:28 AM. MCenter resumed taking beam a bit later.

At 12:56 PM, the MI beam permit tripped off due to beam problems for the first of six times in six and a half hours. It reset each time.

At 2:43 PM, Operations reported that the Pbar target stopped rotating. Operations called an expert who advised them on how to get it moving again.

Operations put the Recycler into shot setup at 4:27 AM. The transfers finished at 5:21 AM.

Pbar resumed stacking and MCenter resumed taking beam at 5:32 AM.

At 3:52 PM, the NuMI beam permit dropped due to beam in the kicker gap for the first of three times in four hours. It reset each time.

At 4:06 PM, the NuMI beam permit tripped off due to beam problems for the first of 12 times in about seven hours and forty minutes. It reset each time.

Operations terminated store 4293 at 7:35 PM and then put the TeV into a dry squeeze.

Operations made the D-Zero (D0) experiment ready for a one hour access at 7:41 PM.

Operations put the TeV into shot setup at 9:28 PM.

At 11:10 PM, a Pbar expert reported that the target had stopped rotating. He said an access would be needed to fix it.

Sunday July 24
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established combination store 4295 at 12:20 AM with an initial luminosity of 99.45E30. Pbar resumed stacking.

MCenter resumed taking beam at 12:37 AM.

Operations established beam to MTest at 1:12 AM.

MTest said that they were through with the beam for the night at 4:08 AM.

Recycler began stashing antiprotons at 4:25 AM.

The stash ended and Pbar resumed stacking at 4:57 AM.

At 5:55 AM, a MI kicker power supply tripped off (I:KPS4A). Operators weren’t able to reset it until 6:40 AM. An expert was called in to investigate.

At 9:18 AM, a Booster RF station (BRF19) tripped off. This was the first of many RF trips of various stations until 5 PM.

Operations put the Recycler into shot setup at 10:59 AM to stash antiprotons.

The stashing ended and Pbar resumed stacking at 11:49 AM.

MCenter resumed taking beam at 11:59 AM.

Operations established beam to MTest at 12:25 PM.

At 6:53 PM, MIRF10 tripped off and took awhile to reset. Operations contacted experts.

At 7:07 PM, the Booster began to suffer multiple RF trips due to station sparking. An expert came in to investigate.

At 7:28 PM, MIRF10 tripped off. While trying to fix MIRF10, all MI beam stopped.

Operators bypassed MIRF10. Pbar resumed stacking and NuMI resumed taking beam at 8:08 PM.

At 8:12 PM, Operations reported that BRF19 was continuing to trip due to sparking.

MTest resumed taking beam at 8:24 PM.

At 8:26 PM, a RF expert arrived. He asked to have the 13.8kV racked out so he could begin MIRF repairs. Only MiniBooNE continued to receive beam.

At 8:36 PM, another RF expert asked to have all the Booster RF stations turned off. This shut all beam off for about fifteen minutes.

The RF expert reported at 9:49 PM that MI was ready to run.

By 9:54 PM beam had returned to Pbar, NuMI, and Switchyard (SY).

At 11:55 PM, Operations reported an emittance blowup in Recycler.

Monday July 25
The midnight shift began with Operations monitoring store 4295 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

At 1:38 AM, NuMI tripped off for the first of a whole bunch of times in about thirty minutes. It reset each time.

MTest finished taking beam for the night at 1:45 AM.

At 2:02 AM, BRF19 began tripping off and taking other RF stations down with it.

Operations terminated store 4295 at 4:05 AM. Operations put the TeV back into shot setup.

At 4:52 AM, BRF19 tripped off the first of many, many times over the next three hours.

Operations established store 4296 at 7:05 AM with an initial luminosity of 116.27E30. Pbar resumed stacking.

MCenter resumed taking beam at 7:27 AM.

The Plan for Monday
The plans for today are to allow a MI and Booster access for RF repairs around 12:30 PM. This will be around a four hour access.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/25/2005   email Fermilab